OpenStudy (anonymous):

Before I ask my question their is a circle at the top that shows 1/3 is Sedan, 1/4 is Van, 1/6 is truck and 1/4 is motorcycle . it says "write a fraction addition problem to show the probablility of spinning a truck or van" i dont get how to do this! please help!

OpenStudy (anonymous):

I think they want you to see this circle as a Twister board (or a casino Roulette). If you were to spin an arrow (or a bottle, like the kissing game) in the center of the circle, the arrow would land on either the "Sedan" section, the "Van" section, the "truck" section or the "motorcycle" section. Now, what you're asked is : what is the probability that the arrow is going to point the Van or the truck section. Do you understand better?

OpenStudy (anonymous):

1/4 + 1/6 = 3/12 +2/12 = 5/12 (I think)

OpenStudy (anonymous):

Kellie's right

OpenStudy (anonymous):